# Project page: https://zhang-tao-whu.github.io/projects/DVIS_DAQ DVIS-DAQ: Improving Video Segmentation via Dynamic Anchor Queries

### v2

@@ Abstract
In summary, our main contributions are:
@@
We have found that current query-based methods fall short in managing objects’ emergence and disappearance due to the considerable transition gap between foreground and background queries. We propose a Dynamic Anchor Queries (DAQ) mechanism to address this challenge. The DAQ mechanism dynamically generates anchor queries for the emergence and disappearance of objects, thereby shortening the gap and effectively overcoming the challenge.
@@ 2 Related Work
Video Instance Segmentation (VIS). Earlier works [70, 39, 71] extend image instance segmentation methods by adding tracking heads and learning the feature association. With the raise of vision transformer [10, 58, 18], current SOTA video instance methods [58, 25, 12, 63, 26, 64, 72, 37, 41, 31, 21] adopt query-based designs. In particular, Video K-Net [38] and IDOL [64] directly learn the query association embeddings via contrastive learning. After that, association-based methods [70, 26, 64, 72, 37, 41], accomplish the tracking of video targets between adjacent frames or adjacent video clips through object matching. Meanwhile, several models [25, 12, 63] directly predict 3D volumes of video instances across times in a semi-online manner. On the other hand, propagation-based methods [24, 31, 21, 75] delegate the tracking problem to the transformer decoder. They take the object queries outputted from the previous frame or video clip as input and iteratively optimize them to segment the objects in the current frame. Note that these methods can also be applied for video panoptic segmentation (VPS) and video semantic segmentation (VSS) tasks [52, 44, 43, 59, 27, 82] by adding queries for segmenting the stuff. However, both poorly handle challenges such as new-emerging objects and target disappearance. Our approach improves the model’s ability to address these challenges by employing dynamic anchor queries to manage the emergence and disappearance of objects.
@@ 3.2 Dynamic Anchor Queries
Dynamic Anchor Queries for Emergence. When processing the T frame of a video, we consider all objects appearing in the current frame as candidates for newly emerging objects. The dynamic anchor queries for emergence DAQ are naturally generated based on the features of these candidates. As shown in Fig. 3 (a), DAQ consists of two parts: query feature DAQ and query positional embedding DAQ. The query feature DAQ is a learnable embedding shared by all dynamic anchor queries. The query positional embedding DAQ is obtained using the appearance features of the candidates through simple Mask-Pooling:
@@
Dynamic Anchor Queries for Disappearance. When processing the T frame, we consider all currently tracked objects as candidates for disappearing objects. Similar to the dynamic anchor queries for emergence, the DAQ for disappearing objects are naturally generated based on the features of these candidates. As shown in Fig. 3 (b), the DAQ also consists of query feature DAQ and query positional embedding DAQ. We use the candidates’ momentum-weighted appearance features as the query positional embedding DAQ. The momentum-weighted function follows CTVIS [72]:
@@
Continuously Tracked Queries. We also assign query positional embedding CTQ for continuously tracked queries (CTQ) to maintain the format consistency with dynamic anchor queries. We use the momentum-weighted appearance feature of the tracked objects as CTQ, the same process as Eq. 3, 4, and 5.
Incorporating Dynamic Anchor Queries into a Tracker. As depicted in Fig. 4, we can seamlessly integrate DAQ into mainstream propagation-based trackers [75, 24], requiring no alterations to the tracker’s network architecture. We instantiate two trackers: Tracker 1, which takes CTQ and DAQ as inputs and is responsible for tracking continuously appearing and newly emerging objects, and Tracker 2, which receives DAQ along with several learnable background embeddings as inputs and identifies disappearing objects in the current frame. It’s important to note that the SoftMax function in Tracker 2 operates along the Q-dimension to prevent multiple dynamic anchor queries from segmenting the same object. Through this pipeline, we can effectively and reasonably manage and model the emergence and disappearance of objects. Despite introducing an additional tracker, it’s worth highlighting that the computational overhead is minimal, constituting less than 2% of the overall computational cost [75], as processing occurs solely at the query level rather than on dense image features.
@@ 3.3 Emergence and Disappearance Simulation
New Emergence Simulation. As shown in Fig. 5 (a), taking the green object as an example, we can simulate it as an emerging object by removing its corresponding query from CTQ. At this point, this green object is not present in CTQ but exists in the current frame, consistent with the state of a real newly emerged object colored in gray. Then, the Tracker needs to transition the corresponding anchor query from DAQ to a foreground query and predict this green object’s segmentation mask and class.
@@ 4.2 Main Experiments
Performance on Youtube-VIS 2022 dataset. Youtube-VIS 2022 has been expanded with 71 long videos based on the 2021 version. Therefore, we only report the AP for the long videos in Tab. 2. When using ResNet-50 as the backbone, DVIS-DAQ outperforms the baseline DVIS with 3.0 AP. Our method outperforms DVIS++ with 4.5 AP (42.0 vs. 37.5) when using VIT-L as the backbone.
@@ 4.3 Ablation Studies and Analysis
We perform ablation experiments on the OVIS dataset using the ResNet-50 backbone and 40K training iterations. We use DVIS [75] as the baseline. When aligning the training and testing settings with DVIS++ [76], the baseline achieves an AP of 35.4. More detailed settings are provided in the supplementary materials.
@@
Qualitative results comparison. In Fig. 7, we present the visual comparison with our baseline. In this challenging scene, there are new-emerging, disappearing, and reappearing video objects. Our method accurately addresses all these scenarios, whereas DVIS suffers from issues including missed new-emerging objects and ID switches. Although DVIS detected object 1 in frame 2, it incorrectly switched the ID of another object to that of object 1. In the case of the reappearance of objects 1 and 2 in Frame 6, our method successfully maintained continuous tracking. In Frame 7, DVIS incorrectly switched the IDs of two nearby disappearing objects to objects 5 and 6, whereas our method correctly handled these disappearing cases.
@@ 0.A.2 Additional Experiments
To further substantiate the efficacy of the Dynamic Attention Query (DAQ), we integrated it into an alternative video segmentation framework, GenVIS [24], culminating in the development of an augmented architecture dubbed GenVIS-DAQ. We employed ResNet-50 as the backbone and conducted experiments on the more challenging OVIS dataset. For a fair comparison, we used GenVIS without the instance prototype memory (IPM) as our baseline and adhered to the original training strategies utilized by GenVIS. The results are reported in Tab. A1. By incorporating DAQ, we achieved an improvement of 1.7 AP for GenVIS. This demonstrates that using DAQ to manage the emergence and disappearance of objects in videos can enhance the capability of query-based video segmentation methods to handle complex scenes.
@@ 0.A.3 More Ablation Studies
Softmax in the disappearance tracker. As depicted in Tab 2(a), compared to the standard cross-attention, which applies SoftMax on the key dimension, we observe that applying SoftMax on the query dimension yields better performance. This is because employing SoftMax on the query dimension prevents different disappearance DAQ from merging the features of the same candidate into the same target query.
Spatio-temporal padding. As shown in Tab. 2(b), prior to the integration of online results into the offline module, it is imperative to apply both temporal and spatial padding. This preparatory step guarantees that the input is appropriately aligned in both time and space to fulfill the requirement of the offline module. Opting to pad with momentum-weighted appearance features rather than zero or learnable padding affords the advantage of tailoring feature information to each individual video object. This tailored padding strategy has been instrumental in achieving a notable improvement in AP (42.7 vs. 38.4). Additionally, performing spatial padding subsequent to temporal padding can further enhance the AP (43.5 vs. 42.7).
@@ 0.A.4 More Qualitative Results
| Method | Backbone | Online | Offline | ||||||||
| AP | AP | AP | AR | AR | AP | AP | AP | AR | AR | ||
| MinVIS [26] | R50 | 25.0 | 45.5 | 24.0 | 13.9 | 29.7 | - | - | - | - | - |
| IDOL [64] | R50 | 28.2 | 51.0 | 28.0 | 14.5 | 38.6 | - | - | - | - | - |
| GRAtt-VIS [21] | R50 | 36.2 | 60.8 | 36.8 | 16.8 | 40.0 | - | - | - | - | - |
| TCOVIS [31] | R50 | 35.3 | 60.7 | 36.6 | 15.7 | 39.5 | - | - | - | - | - |
| UNINEXT [68] | R50 | 34.0 | 55.5 | 35.6 | - | - | - | - | - | - | - |
| CTVIS [72] | R50 | 35.5 | 60.8 | 34.9 | 16.1 | 41.9 | - | - | - | - | - |
| RefineVIS [1] | R50 | 33.7 | 56.2 | 34.8 | 15.6 | 39.8 | - | - | - | - | - |
| VITA [25] | R50 | - | - | - | - | - | 19.6 | 41.2 | 17.4 | 11.7 | 26.0 |
| GenVIS [24] | R50 | 35.8 | 60.8 | 36.2 | 16.3 | 39.6 | 34.5 | 59.4 | 35.0 | 16.6 | 38.3 |
| DVIS [75] | R50 | 30.2 | 55.0 | 30.5 | 14.5 | 37.3 | 33.8 | 60.4 | 33.5 | 15.3 | 39.5 |
| DVIS++ [76] | R50 | 37.2 | 62.8 | 37.3 | 15.8 | 42.9 | 41.2 | 68.9 | 40.9 | 16.8 | 47.3 |
| DVIS-DAQ | R50 | 38.7 | 65.5 | 37.6 | 17.0 | 45.2 | 43.5 | 72.5 | 43.8 | 16.4 | 49.1 |
| MinVIS [26] | Swin-L | 39.4 | 61.5 | 41.3 | 18.1 | 43.3 | - | - | - | - | - |
| IDOL [64] | Swin-L | 40.0 | 63.1 | 40.5 | 17.6 | 46.4 | - | - | - | - | - |
| GRAtt-VIS [21] | Swin-L | 45.7 | 69.1 | 47.8 | 19.2 | 49.4 | - | - | - | - | - |
| TCOVIS [31] | Swin-L | 46.7 | 70.9 | 49.5 | 19.1 | 50.8 | - | - | - | - | - |
| RefineVIS [1] | Swin-L | 46.1 | 69.7 | 47.8 | 19.0 | 50.8 | - | - | - | - | - |
| UNINEXT [68] | VIT-H | 49.0 | 72.5 | 52.2 | - | - | - | - | - | - | - |
| CTVIS [72] | Swin-L | 46.9 | 71.5 | 47.5 | 19.1 | 52.1 | - | - | - | - | - |
| VITA [25] | Swin-L | - | - | - | - | - | 27.7 | 51.9 | 24.9 | 14.9 | 33.0 |
| GenVIS [24] | Swin-L | 45.2 | 69.1 | 48.4 | 19.1 | 48.6 | 45.4 | 69.2 | 47.8 | 18.9 | 49.0 |
| DVIS [75] | Swin-L | 45.9 | 71.1 | 48.3 | 18.5 | 51.5 | 48.6 | 74.7 | 50.5 | 18.8 | 53.8 |
| DVIS++ [76] | VIT-L | 49.6 | 72.5 | 55.0 | 20.8 | 54.6 | 53.4 | 78.9 | 58.5 | 21.1 | 58.7 |
| DVIS-DAQ | Swin-L | 49.5 | 76.1 | 51.7 | 19.5 | 54.9 | - | - | - | - | - |
| DVIS-DAQ | VIT-L | 53.7 | 78.7 | 58.2 | 20.8 | 59.5 | 57.1 | 83.8 | 62.9 | 21.0 | 62.3 |
| Method | Backbone | Youtube-VIS 2019 | Youtube-VIS 2021 | ||||||||
| AP | AP | AP | AR | AR | AP | AP | AP | AR | AR | ||
| Online | |||||||||||
| SipMask [9] | R50 | 33.7 | 54.1 | 35.8 | 35.4 | 40.1 | 31.7 | 52.5 | 34.0 | 30.8 | 37.8 |
| CrossVIS [71] | R50 | 36.3 | 56.8 | 38.9 | 35.6 | 40.7 | 34.2 | 54.4 | 37.9 | 30.4 | 38.2 |
| VISOLO [20] | R50 | 38.6 | 56.3 | 43.7 | 35.7 | 42.5 | 36.9 | 54.7 | 40.2 | 30.6 | 40.9 |
| MinVIS [26] | R50 | 47.4 | 69.0 | 52.1 | 45.7 | 55.7 | 44.2 | 66.0 | 48.1 | 39.2 | 51.7 |
| IDOL [64] | R50 | 49.5 | 74.0 | 52.9 | 47.7 | 58.7 | 43.9 | 68.0 | 49.6 | 38.0 | 50.9 |
| GenVIS [24] | R50 | 50.0 | 71.5 | 54.6 | 49.5 | 59.7 | 47.1 | 67.5 | 51.5 | 41.6 | 54.7 |
| GRAtt-VIS [21] | R50 | 50.4 | 70.7 | 55.2 | 48.4 | 58.7 | 48.9 | 69.2 | 53.1 | 41.8 | 56.0 |
| TCOVIS [31] | R50 | 52.3 | 73.5 | 57.6 | 49.8 | 60.2 | 49.5 | 71.2 | 53.8 | 41.3 | 55.9 |
| UNINEXT [68] | R50 | 53.0 | 75.2 | 59.1 | - | - | - | - | - | - | - |
| DVIS [75] | R50 | 51.2 | 73.8 | 57.1 | 47.2 | 59.3 | 46.4 | 68.4 | 49.6 | 39.7 | 53.5 |
| DVIS++ [76] | R50 | 55.5 | 80.2 | 60.1 | 51.1 | 62.6 | 50.0 | 72.2 | 54.5 | 42.8 | 58.4 |
| DVIS-DAQ | R50 | 55.2 | 78.7 | 61.9 | 50.6 | 63.7 | 50.4 | 72.4 | 55.0 | 41.8 | 57.6 |
| MinVIS [26] | Swin-L | 61.6 | 83.3 | 68.6 | 54.8 | 66.6 | 55.3 | 76.6 | 62.0 | 45.9 | 60.8 |
| IDOL [64] | Swin-L | 64.3 | 87.5 | 71.0 | 55.6 | 69.1 | 56.1 | 80.8 | 63.5 | 45.0 | 60.1 |
| GenVIS [24] | Swin-L | 64.0 | 84.9 | 68.3 | 56.1 | 69.4 | 59.6 | 80.9 | 65.8 | 48.7 | 65.0 |
| GRAtt-VIS [21] | Swin-L | 63.1 | 85.6 | 67.2 | 55.5 | 67.8 | 60.3 | 81.3 | 67.1 | 48.8 | 64.5 |
| RefineVIS [1] | Swin-L | 64.3 | 87.6 | 70.9 | 55.8 | 68.2 | 61.4 | 84.1 | 68.5 | 48.3 | 65.2 |
| CTVIS [72] | Swin-L | 65.6 | 87.7 | 72.2 | 56.5 | 70.4 | 61.2 | 84.0 | 68.8 | 48.0 | 65.8 |
| TCOVIS [31] | Swin-L | 64.1 | 86.6 | 69.5 | 55.8 | 69.0 | 61.3 | 82.9 | 68.0 | 48.6 | 65.1 |
| UNINEXT [68] | VIT-H | 66.9 | 87.5 | 75.1 | - | - | - | - | - | - | - |
| DVIS [75] | Swin-L | 63.9 | 87.2 | 70.4 | 56.2 | 69.0 | 58.7 | 80.4 | 66.6 | 47.5 | 64.6 |
| DVIS++ [76] | VIT-L | 67.7 | 88.8 | 75.3 | 57.9 | 73.7 | 62.3 | 82.7 | 70.2 | 49.5 | 68.0 |
| DVIS-DAQ | Swin-L | 65.7 | 88.1 | 73.6 | 55.6 | 70.7 | 61.1 | 84.2 | 68.2 | 48.0 | 66.6 |
| DVIS-DAQ | VIT-L | 68.3 | 88.5 | 76.1 | 58.0 | 73.5 | 62.4 | 83.6 | 70.8 | 49.1 | 68.0 |
| Offline | |||||||||||
| SeqFormer [63] | Swin-L | 59.3 | 82.1 | 66.4 | 51.7 | 64.4 | 51.8 | 74.6 | 58.2 | 42.8 | 58.1 |
| M2F-VIS [12] | Swin-L | 60.4 | 84.4 | 67.0 | - | - | 52.6 | 76.4 | 57.2 | - | - |
| VITA [25] | Swin-L | 63.0 | 86.9 | 67.9 | 56.3 | 68.1 | 57.5 | 80.6 | 61.0 | 47.7 | 62.6 |
| GenVIS [24] | Swin-L | 63.8 | 85.7 | 68.5 | 56.3 | 68.4 | 60.1 | 80.9 | 66.5 | 49.1 | 64.7 |
| MDQE [32] | Swin-L | 59.9 | - | - | - | - | 56.2 | 80.0 | 61.1 | 44.9 | 59.1 |
| NOVIS [41] | Swin-L | 65.7 | 87.8 | 72.2 | 56.3 | 70.3 | 59.8 | 82.0 | 66.5 | 47.9 | 64.6 |
| Tube-Link [37] | Swin-L | 64.6 | 86.6 | 71.3 | 55.9 | 69.1 | 58.4 | 79.4 | 64.3 | 47.5 | 63.6 |
| RefineVIS [1] | Swin-L | 64.4 | 88.3 | 72.2 | 55.8 | 68.4 | 61.2 | 83.7 | 69.2 | 47.9 | 64.8 |
| DVIS [75] | Swin-L | 64.9 | 88.0 | 72.7 | 56.5 | 70.3 | 60.1 | 83.0 | 68.4 | 47.7 | 65.7 |
| DVIS++ | VIT-L | 68.3 | 90.3 | 76.1 | 57.8 | 73.4 | 63.9 | 86.7 | 71.5 | 48.8 | 69.5 |
| DVIS-DAQ | VIT-L | 69.2 | 90.8 | 76.8 | 58.4 | 75.5 | 64.5 | 86.3 | 72.4 | 49.6 | 70.7 |

### v4

@@ Abstract
In summary, our main contributions are as follows:
@@
We have found that current query-based methods fall short in managing objects’ emergence and disappearance due to the considerable transition gap between foreground and background queries. We propose a Dynamic Anchor Queries (DAQ) mechanism to address this challenge. The DAQ mechanism dynamically generates anchor queries for objects’ emergence and disappearance, thereby shortening the gap and effectively overcoming the challenge.
@@ 2 Related Work
Video Instance Segmentation (VIS). Earlier works [66, 35, 67] extend image instance segmentation methods by adding tracking heads and learning the feature association. With the rise of vision transformer [8, 54, 16], current SOTA video instance methods [54, 22, 10, 59, 23, 60, 68, 33, 37, 28, 18] adopt query-based designs. In particular, Video K-Net [34] and IDOL [60] directly learn the query association embeddings via contrastive learning. After that, association-based methods [66, 23, 60, 68, 33, 37], accomplish the tracking of video targets between adjacent frames or adjacent video clips through object matching. Meanwhile, several models [22, 10, 59] directly predict 3D volumes of video instances across times in a semi-online manner. On the other hand, propagation-based methods [21, 28, 18, 71] delegate the tracking problem to the transformer decoder. They take the object queries outputted from the previous frame or video clip as input and iteratively optimize them to segment the objects in the current frame. Note that these methods can also be applied for video panoptic segmentation (VPS) and video semantic segmentation (VSS) tasks [48, 40, 39, 55, 24, 78] by adding queries for segmenting the stuff. However, both poorly handle challenges such as new-emerging objects and target disappearance. Our approach improves the model’s ability to address these challenges by employing dynamic anchor queries to manage the emergence and disappearance of objects.
@@ 3.2 Dynamic Anchor Queries
Dynamic Anchor Queries for Emergence. When processing the Tth frame of a video, we consider all objects appearing in the current frame as candidates for newly emerging objects. The dynamic anchor queries for emergence DAQEmg are naturally generated based on the features of these candidates. As shown in Fig. 3 (a), DAQEmg consists of two parts: query feature DAQ and query positional embedding DAQ. The query feature DAQ is a learnable embedding shared by all dynamic anchor queries. The query positional embedding DAQ is obtained using the appearance features of the candidates through simple Mask-Pooling:
@@
Dynamic Anchor Queries for Disappearance. When processing the Tth frame, we consider all currently tracked objects as candidates for disappearing objects. Similar to the dynamic anchor queries for emergence, the DAQDis for disappearing objects are naturally generated based on the features of these candidates. As shown in Fig. 3 (b), the DAQDis also consists of query feature DAQ and query positional embedding DAQ. We use the candidates’ momentum-weighted appearance features as the query positional embedding DAQ. The momentum-weighted function follows CTVIS [68]:
@@
Continuously Tracked Queries. We also assign query positional embedding CTQpos for continuously tracked queries (CTQ) to maintain the format consistency with dynamic anchor queries. We use the momentum-weighted appearance feature of the tracked objects as CTQpos, the same process as Eq. 3, 4, and 5.
Incorporating Dynamic Anchor Queries into a Tracker. As depicted in Fig. 4, we can seamlessly integrate DAQ into mainstream propagation-based trackers [71, 21], requiring no alterations to the tracker’s network architecture. We instantiate two trackers: Tracker 1, which takes CTQ and DAQEmg as inputs and is responsible for tracking continuously appearing and newly emerging objects, and Tracker 2, which receives DAQDis along with several learnable background embeddings as inputs and identifies disappearing objects in the current frame. It’s important to note that the SoftMax function in Tracker 2 operates along the Q-dimension to prevent multiple dynamic anchor queries from segmenting the same object. Through this pipeline, we can effectively and reasonably manage and model the emergence and disappearance of objects. Despite introducing an additional tracker, it’s worth highlighting that the computational overhead is minimal, constituting less than 2% of the overall computational cost [71], as processing occurs solely at the query level rather than on dense image features.
@@ 3.3 Emergence and Disappearance Simulation
New Emergence Simulation. As shown in Fig. 5 (a), taking the green object as an example, we can simulate it as an emerging object by removing its corresponding query from CTQ. At this point, this green object is not present in CTQ but exists in the current frame, consistent with the state of a real newly emerged object colored in gray. Then, the Tracker needs to transition the corresponding anchor query from DAQEmg to a foreground query and predict this green object’s segmentation mask and class.
@@ 4.2 Main Experiments
Performance on Youtube-VIS 2022 dataset. Youtube-VIS 2022 has been expanded with 71 long videos based on the 2021 version. Therefore, we only report the APL for the long videos in Tab. 2. When using ResNet-50 as the backbone, DVIS-DAQ outperforms the baseline DVIS with 3.0 APL. Our method outperforms DVIS++ with 4.5 APL (42.0 vs. 37.5) when using VIT-L as the backbone.
@@ 4.3 Ablation Studies and Analysis
We perform ablation experiments on the OVIS dataset using the ResNet-50 backbone and 40K training iterations. We use DVIS [71] as the baseline. When aligning the training and testing settings with DVIS++ [72], the baseline achieves an AP of 35.4. More detailed settings are provided in the supplementary materials. Moreover, we also refer the readers for more different architectures using our DAQ in supplementary part.
@@
Qualitative results comparison. In Fig. 7, we present the visual comparison with our baseline. In this challenging scene, there are new-emerging, disappearing, and reappearing video objects. Our method accurately addresses all these scenarios, whereas DVIS suffers from issues including missed new-emerging objects and ID switches. Although DVIS detected object 1 in frame 2, it incorrectly switched the ID of another object to that of object 1. Regarding the reappearance of objects 1 and 2 in Frame 6, our method successfully maintained continuous tracking. In Frame 7, DVIS incorrectly switched the IDs of two nearby disappearing objects to objects 5 and 6, whereas our method correctly handled these disappearing cases.
@@ 0.A.2 Additional Experiments
Generalization ability of DAQ. To further substantiate the efficacy of the Dynamic Attention Query (DAQ), we integrated it into an alternative video segmentation framework, GenVIS [21], culminating in the development of an augmented architecture dubbed GenVIS-DAQ. We employed ResNet-50 as the backbone and conducted experiments on the more challenging OVIS dataset. For a fair comparison, we used GenVIS without the instance prototype memory (IPM) as our baseline and adhered to the original training strategies utilized by GenVIS. The results are reported in Tab. A1. By incorporating DAQ, we achieved an improvement of 1.7 AP for GenVIS. This demonstrates that using DAQ to manage the emergence and disappearance of objects in videos can enhance the capability of query-based video segmentation methods to handle complex scenes.
Performance improvements on emergence and disappearance. We conduct separate evaluations for newly emerged and disappeared objects on the OVIS and VIPSeg datasets, with the results presented in Tab .A2. For newly emerged and disappeared objects, our method outperforms the baseline DVIS [71] by 7.3 VPQ and 6.8 AP on VIPSeg and OVIS, respectively, indicating the effectiveness of our method for this problem.
Feature transition gap. We claimed that our DAQ mechanism enhances the management of objects’ emergence and disappearance by shortening the feature transition gap between the anchor query and the target. We define the transition gap as the feature distance from the anchor query to the target. In DVIS [71], the anchor query for a newly emerged object is a background query, and the target is the feature representation of the object. In contrast, the anchor query for a disappeared object is the feature representation of the object before it disappeared, and the target is a background feature. In our method, the dynamic anchor query is a mixture of background and foreground for both situations, where learnable embeddings and candidates provide the background and foreground features, respectively. Our DAQs offer information on potential future states of targets. Therefore, our method results in a smaller transition gap to the target. As shown in Tab. A3, we calculate the transition gap for DVIS and our method, where our DAQ significantly shortens the transition gap.
@@
| Method | VIPSeg | OVIS | ||
| VPQall | VPQed | APall | APed | |
| DVIS [71] | 38.7 | 29.0 | 26.9 | 19.1 |
| DVIS+DAQ | 42.1 | 36.3 | 30.2 | 25.9 |
| Improvement | +3.4 | +7.3 | +3.3 | +6.8 |
| Methods | CS | NED |
| DVIS [71] | ||
| DVIS+DAQ |
@@ 0.A.3 More Ablation Studies
Softmax in the disappearance tracker. As depicted in Tab. 4(a), compared to the standard cross-attention, which applies SoftMax on the key dimension, we observe that applying SoftMax on the query dimension yields better performance. This is because employing SoftMax on the query dimension prevents different disappearance DAQ from merging the features of the same candidate into the same target query.
Spatio-temporal padding. As shown in Tab. 4(c), before integrating online results into the offline module, it is imperative to apply both temporal and spatial padding. This preparatory step guarantees that the input is appropriately aligned in both time and space to fulfill the requirement of the offline module. Opting to pad with momentum-weighted appearance features rather than zero or learnable padding affords the advantage of tailoring feature information to each individual video object. This tailored padding strategy has been instrumental in achieving a notable improvement in AP (42.7 vs. 38.4). Additionally, performing spatial padding subsequent to temporal padding can further enhance the AP (43.5 vs. 42.7).
Computation cost analysis. The computational cost of DVIS-DAQ components was measured by evaluating the parameters, FLOPs, and inference time of the segmenter and tracker with DAQ. Tab. As shown in Tab. 4(b), the additional tracker #2 only accounts for 1.6% of the total computational cost.
@@
| Component | Params (M) | FLOPs (G) | Time (ms) |
| Segmenter | 44.15 | 225.14 | 50.01 |
| Tracker #1 | 9.94 | 3.78 | 10.32 |
| Tracker #2 | 8.68 | 3.67 |